OpenStudy (yolo_king):

A bag contains 3 red marbles, 2 blue marbles and 5 green marbles. What is the probability of selecting a blue marble, replacing it in the bag and then selecting a green marble? A. 7/10 B. 4/5 C. 3/20 D. 1/10

OpenStudy (yolo_king):

I got 1/10 can someone confirm?

OpenStudy (moldybubblegum12):

If you replace the marble inside the bag, then the total number of marbles remains the same. (2/10) * (5/10) ==> 10/100

OpenStudy (yolo_king):

1/10 * 1/10 = 1/10

OpenStudy (yolo_king):

Thanks!

OpenStudy (moldybubblegum12):

\[\frac{ 10 }{ 100 }\] is the same as \[\frac{ 1 }{ 10 }\]